### v1

@@ Abstract
Context It is commonly accepted that the quality of requirements specifications impacts subsequent software engineering activities. However, we still lack empirical evidence to support organizations in deciding whether their requirements are good enough or impede subsequent activities. Objective We aim to contribute empirical evidence to the effect that requirements quality defects have on a software engineering activity that depends on this requirement. Method We replicate a controlled experiment in which 25 participants from industry and university generate domain models from four natural language requirements containing different quality defects. We evaluate the resulting models using both frequentist and Bayesian data analysis. Results Contrary to our expectations, our results show that the use of passive voice only has a minor impact on the resulting domain models. The use of ambiguous pronouns, however, shows a strong effect on various properties of the resulting domain models. Most notably, ambiguous pronouns lead to incorrect associations in domain models. Conclusion Despite being equally advised against by literature and frequentist methods, the Bayesian data analysis shows that the two investigated quality defects have vastly different impacts on software engineering activities and, hence, deserve different levels of attention. Our employed method can be further utilized by researchers to improve reliable, detailed empirical evidence on requirements quality.
@@ 1 Introduction
Software requirements specify the needs and constraints stakeholders impose on a desired system. Software requirements specifications (SRS), the explicit manifestation of requirements as an artifact mendez2019artefacts , serve as input for various subsequent software engineering (SE) activities, such as deriving a software architecture, implementing features, or generating test cases mendez2015artefact . As a consequence, the quality of an SRS impacts the quality of requirements-dependent activities frattini2023requirements ; femmer2018requirements ; femmer2015ABREQM . A quality defect in an SRS—for example, an ambiguous formulation—can cause differing interpretations and result in the design and implementation of a solution that does not meet the stakeholders’ needs mendez2017naming . The inherent complexity of natural language (NL), which is most commonly used for specifying requirements franch2017practitioners , aggravates this challenge further. Since quality defects are understood to scale in cost for removal boehm1984software , organizations are interested in identifying and removing these defects as early as possible montgomery2022empirical .
@@
In this article, we aim to address the above-mentioned shortcomings by i) replicating the only controlled experiment known to us that contributes empirical evidence on the impact of passive voice on requirements quality femmer2014impact , and strengthening the robustness of this evidence. Further, we ii) collect data about relevant context factors and integrate these data in our data analysis. Finally, we iii) contrast the state-of-the-art frequentist data analysis (FDA) with a Bayesian data analysis (BDA). The latter has recently been popularized in SE research furia2019bayesian since it generates more nuanced empirical insights. Our study is categorized as a laboratory experiment in a contrived setting stol2018abc , isolating the effect of the selected quality factors of interest. The causal inference of their impact contributes to our long-term goal of providing an empirically grounded understanding of the impact of requirements quality. This will support organizations in assessing their requirements and detecting relevant quality defects early.
To that end, we conducted an experiment with 25 participants simulating a requirements-dependent activity (i.e., domain modeling) using four natural-language requirements as input. These requirements contained two common quality defects, passive voice femmer2014impact and ambiguous pronouns ezzini2022automated . Our experiment includes context factors such as experience in software engineering (SE) and RE, domain knowledge, and task experience.
@@
    a replication of the only controlled experiment investigating the impact of requirements quality;
@@
    the application of BDA to requirements quality research, which is among the first of its kind in RE.
The remainder of this manuscript is organized as follows. Section 2 introduces relevant related work. We present our research method in Section 3 and the results in Section 4. We discuss these results in Section 5 before concluding our manuscript in Section 6. All supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results, are available in our replication package [^1] The replication package is available at https://doi.org/10.5281/zenodo.10423666.
@@ 2.1 Requirements Quality
2.1.1 Requirements Quality in General
@@
Requirements quality research seeks answers to this need montgomery2022empirical . One main driver of this research is requirements quality factors frattini2022live , i.e., metrics that can be evaluated on NL requirements specifications to determine quality defects. For example, the voice of a NL sentence (active or passive) is considered a quality factor, as the use of passive voice is associated with bad requirements quality due to potential omission of information femmer2014impact . Automatic detection techniques using natural language processing (NLP) zhao2021natural can automatically evaluate quality factors to detect defects in NL requirements specifications femmer2017rapid .
@@ 2.2 Bayesian data analysis in software engineering
In recent years, SE research has adopted Bayesian data analysis (BDA) for statistical evaluations. BDA signifies a departure from frequentist methods like null-hypothesis significance testing (NHST), the previous state-of-the-art in terms of inferential statistics in SE research. NHST determines whether there is a “statistically significant” difference between two or more distributions. Observations of a dependent variable are stratified by an independent variable to obtain a binary answer of whether or not different values of the independent variable correlate with different distributions of the dependent variable.
Opposed to that, BDA abstains from reducing complex variable distributions to binary inference mcelreath2020statistical . Instead, dependent variables are expressed as a probability distribution, which preserves the natural uncertainty with which any variable is determined. Similarly, the impact of any independent variable on the dependent variable is expressed in terms of a probability distribution. Using Bayes’ Theorem, these assigned prior probability distributions are updated with observed data to obtain a posterior probability distribution furia2019bayesian . Given the observed data, these posterior probability distributions model the most likely impact of variable values. BDA methods are becoming widely adopted also due to the modern computational power enabling Markov Chain Monte Carlo (MCMC) randomized algorithms brooks2011handbook , tools like Stan carpenter2017stan , and libraries like rethinking mcelreath2020statistical and brms burkner2017brms .
@@
    BDA makes causal assumptions explicit. The Bayesian workflow gelman2020bayesian makes any hypothesis of causal relations explicit. Analyses become more transparent, and competing causal assumptions are easier to assess.
@@ 3 Method
We replicate a controlled experiment that investigates the impact of requirements quality on a software engineering activity. Our goal is both to (1) contribute empirical evidence to the effect of quality defects and (2) compare the inferential capabilities of frequentist (FDA) with Bayesian (BDA) statistics. Our experiment is an external, differentiated replication baldassarre2014replication of the study conducted and reported by Femmer et al. femmer2014impact . In this section, we report the design of the replication following the guidelines by Carver carver2010towards .
@@ 3.1 Goals
RQ1 is dedicated to the main relationship of interest between quality defects and an affected activity. RQ1.1 directly maps onto the general research question driving the original study femmer2014impact . RQ1.2 extends the scope of the investigation of quality factors with ambiguous pronouns. RQ1.3 investigates the interaction between the two quality factors. RQ2 adds a context-sensitive perspective to the relationship. RQ2.1 focuses on the direct effect that context factors have on the affected activity. RQ2.2 investigates whether context factors mediate the effect of quality defects on the activity.
@@ 3.2 Original Experiment
For the dependent variable, the authors calculated the number of missing domain model elements (i.e., actors, objects, and associations). Although the authors also recorded context variables, such as a categorical assessment of general knowledge in SE and RE, these were not used in the analysis. The analysis followed a frequentist approach performing a null-hypothesis significance test for each of the three domain model elements to determine whether a statistically significant difference between the experimental groups exists. The study shows a statistically significant difference in the number of identified associations but not in the number of actors or objects. The authors conclude that the commonly assumed impact of passive voice on missing domain model actors is actually negligible, but passive voice impedes the understanding of the relationships between entities in the requirements specification.
@@ 3.3 Our Experiment
We replicate the study by Femmer et al. femmer2014impact because it is the only study that contributes empirical evidence to the impact of passive voice on requirements quality montgomery2022empirical ; frattini2022live in a controlled, experimental setting. During the preparation of our study, we conferred with the authors of the original study. Through this interaction, we identified several opportunities to improve the experimental design. Consequently, we made the following changes to the original study.
@@
  • Dependent variables: We merged two types of elements in the domain model (the nodes of the model, i.e., actors and objects) into a single type entity. This decision was taken after feedback from a pilot study. Furthermore, we increased the dependent variables by additionally evaluating the number of superfluous entities, the number of wrong associations, and the duration for creating the domain model.
@@
  • Analysis: The crossover design produces paired data as opposed to the unpaired data of the original experiment, which changes the appropriate hypothesis test vegas2015crossover (t-test in the original study vs. Wilcoxon rank sum test in this study). In addition, we extend the original FDA by performing a Bonferroni correction to deal with family-wise error rate when testing multiple hypotheses benjamini1995controlling . Furthermore, we additionally analyze the data using BDA.
In the rest of this subsection, we report the design of our experiment following the guidelines by Jedlitschka et al. jedlitschka2008reporting .
@@ 3.3.2 Hypotheses
[…]
@@
Requirements Quality Defect
RQD
ind
The use of a verb in passive voice, an ambiguous pronoun, or both
categorical
{none, PV, AP, PVAP}
Experience in SE
exp.se
con
Years of experience in software engineering
count
$\mathbb{N}$
Experience in RE
exp.re
con
Years of experience in requirements engineering
count
$\mathbb{N}$
Education
edu
con
Highest acquired degree
ordinal
{High School, B.Sc., M.Sc., Ph.D.}
Primary role
role
con
SE-related Role with the most years of professional experience
categorical
{requirements engineer, product owner, software architect, developer, tester, quality engineer, trainer, manager, none}
Task experience
exp.task
con
Experience with the task of domain modeling
ordinal
{never, rarely, from time to time, often}
Formal modeling training
formal
con
Formal training in in domain modeling
categorical
{true,false}
Domain knowledge in {domain}
dom. {domain}
con
@@
ordinal
{1,2,3,4,5}
Duration
$D$
dep
Number of minutes it took the participant to complete the experimental task on one requirement
count
$\mathbb{N}$
Missing entities
$E^{-}$
dep
Number of relevant entities missing from the submitted domain model
count
$[0,E_{expected}]$
Superfluous entities
$E^{+}$
dep
Number of not relevant entities included in the submitted domain model
count
$\mathbb{N}$
Missing associations
$A^{-}$
dep
Number of associations missing from the submitted domain model
count
$[0,A_{expected}]$
Wrong associations
$A^{\times}$
dep
Number of associations where either the source or target of the edge is a different entity than implied by the requirement
count
$[0,A_{found}]$
@@ 3.3.3 Experimental Design
Given our sampling strategy involving industry practitioners, which are difficult to recruit for controlled experiments sjoberg2003challenges , we anticipated a moderate sample size of participants. Consequently, we opted for a crossover design instead of a parallel design, i.e., we apply every treatment to all subjects instead of distributing the subjects among the treatments. Previously, Kitchenham et al. advised against the use of crossover designs kitchenham2003case . Mainly, the validity of crossover design experiments is challenged by the following confounding factors:
@@
    the period in which a treatment is applied to a subject, as certain periods may influence the dependent variables (e.g., measuring the dependent variable in the period right before an exam, participating students may be more tense and less focused);
@@
    the sequence, in which the treatments are applied to a subject, as certain sequences may entail a learning effect that systematically influences a dependent variable (e.g., repeatedly performing a task might improve the performance of experiment participants); and
@@
    the previous treatment (also called the carryover effect, as the application of a treatment may have an effect on the dependent variable when applying a subsequent treatment vegas2015crossover .
However, recent adoptions of best practices from other disciplines made this design applicable to SE research without compromising the validity of the results vegas2015crossover ; fucci2016an . The threats to validity can be mitigated during design and analysis vegas2015crossover by (1) randomizing the order of treatments and (2) including the independent variables’ period, sequence, and the interaction between them in the analysis. When controlling the threats to validity, the crossover design provides two benefits. Firstly, it requires fewer participants, as an experiment with $n_{p}$ participants and $n_{t}$ treatments yields $n_{p}\times n_{t}$ observations instead of only $n_{p}$ wohlin2012experimentation . Secondly, it accounts for differences between participants, as the dependent variables can be measured in relation to the average response of each subject instead of the average response of each treatment group kitchenham2003case . Therefore, each subject acts as its own control and mitigates within-subject variance.

### v3

@@ Abstract
It is commonly accepted that the quality of requirements specifications impacts subsequent software engineering activities. However, we still lack empirical evidence to support organizations in deciding whether their requirements are good enough or impede subsequent activities. We aim to contribute empirical evidence to the effect that requirements quality defects have on a software engineering activity that depends on this requirement. We conduct a controlled experiment in which 25 participants from industry and university generate domain models from four natural language requirements containing different quality defects. We evaluate the resulting models using both frequentist and Bayesian data analysis. Contrary to our expectations, our results show that the use of passive voice only has a minor impact on the resulting domain models. The use of ambiguous pronouns, however, shows a strong effect on various properties of the resulting domain models. Most notably, ambiguous pronouns lead to incorrect associations in domain models. Despite being equally advised against by literature and frequentist methods, the Bayesian data analysis shows that the two investigated quality defects have vastly different impacts on software engineering activities and, hence, deserve different levels of attention. Our employed method can be further utilized by researchers to improve reliable, detailed empirical evidence on requirements quality.
@@ 1 Introduction
Software requirements specify the needs and constraints that stakeholders impose on a desired system. Software requirements specifications (SRS), the explicit manifestation of requirements as an artifact mendez2019artefacts , serve as input for various subsequent software engineering (SE) activities, such as deriving a software architecture, implementing features, or generating test cases mendez2015artefact . As a consequence, the quality of an SRS impacts the quality of requirements-dependent activities frattini2023requirements ; femmer2018requirements ; femmer2015ABREQM . A quality defect in an SRS—for example, an ambiguous formulation—can cause differing interpretations and result in the design and implementation of a solution that does not meet the stakeholders’ needs mendez2017naming . The inherent complexity of natural language (NL), which is most commonly used for specifying requirements franch2017practitioners , aggravates this challenge further. Since quality defects are understood to scale in cost for removal boehm1984software , organizations are interested in identifying and removing these defects as early as possible montgomery2022empirical .
@@
In this article, we aim to address the above-mentioned shortcomings by i) conducting a controlled experiment with 25 participants simulating a requirements-dependent activity (i.e., domain modeling) using four natural-language requirements as input. The experiment contributes empirical evidence on the impact of two commonly researched quality factors passive voice femmer2014impact and ambiguous pronouns ezzini2022automated . The investigation of the impact of passive voice is a conceptual replication baldassarre2014replication of the only controlled experiment studying the impact of passive voice on domain modeling femmer2014impact known to us. Therefore, our experiment also strengthens the robustness of their conclusions by providing diagnostic evidence nosek2020replication . Further, we ii) collect data about relevant context factors such as experience in software engineering (SE) and RE, domain knowledge, and task experience, and integrate these data in our data analysis. Finally, we iii) contrast the state-of-the-art frequentist data analysis (FDA) with Bayesian data analysis (BDA), which entails both a causal framework and Bayesian modeling for statistical causal inference mcelreath2020statistical . The latter has recently been popularized in SE research furia2019bayesian since it generates more nuanced empirical insights. Our study is categorized as a laboratory experiment in a contrived setting stol2018abc , isolating the effect of the selected quality factors of interest. The causal inference of their impact contributes to our long-term goal of providing an empirically grounded understanding of the impact of requirements quality. This will support organizations in assessing their requirements and detecting relevant quality defects early.
@@
    a controlled experiment investigating the impact of requirements quality;
@@
    a conceptual replication of the only controlled experiment investigating the impact of passive voice femmer2014impact ;
  3. 3.
    the application of BDA to requirements quality research, which is among the first of its kind in RE; and
  4. 4.
    an archived replication package containing all supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results frattini2024replication .
The remainder of this manuscript is organized as follows. Section 2 introduces relevant related work. We present our research method in Section 3 and the results in Section 4. We discuss these results in Section 5 before concluding our manuscript in Section 6.
@@ 2.1 Requirements Quality
2.1.1 Requirements Quality Research
@@
Requirements quality research seeks answers to this need montgomery2022empirical . One main driver of this research is requirements quality factors frattini2022live , i.e., metrics that can be evaluated on NL requirements specifications to determine quality defects. For example, the voice of an NL sentence (active or passive) is considered a quality factor, as the use of passive voice is associated with bad requirements quality due to potential omission of information femmer2014impact . Automatic detection techniques using natural language processing (NLP) zhao2021natural can automatically evaluate quality factors to detect defects in NL requirements specifications femmer2017rapid .
@@ 2.1.3 Shortcomings in Requirements Quality Research
Requirements Quality Research Gaps Requirements quality research suffers from (1) a lack of empirical evidence about the relevance of quality factors, (2) a lack of context-sensitivity, and (3) evaluations of impact that are more fine-grained than categorical.
@@ 2.2 Bayesian data analysis in software engineering
In recent years, SE research has adopted Bayesian data analysis (BDA) for statistical causal inference. BDA signifies a departure from frequentist methods like null-hypothesis significance testing (NHST), the previous state-of-the-art in terms of inferential statistics in SE research. NHST determines whether there is a “statistically significant” difference between two or more distributions. Observations of a dependent variable are stratified by an independent variable to obtain a binary answer of whether or not different values of the independent variable correlate with different distributions of the dependent variable.
Opposed to that, BDA encourages the use of causal frameworks mcelreath2020statistical . These frameworks make causal assumptions explicit elwert2013graphical and allow reasoning about causally relevant variables pearl1995bayesian ; pearl2016causal . Furthermore, BDA abstains from reducing complex variable distributions to binary inference mcelreath2020statistical . Instead, dependent variables are expressed as a probability distribution, which preserves the natural uncertainty with which any variable is determined. Similarly, the impact of any independent variable on the dependent variable is expressed in terms of a probability distribution. Using Bayes’ Theorem, these assigned prior probability distributions are updated with observed data to obtain a posterior probability distribution furia2019bayesian . Given the observed data, these posterior probability distributions model the most likely impact of variable values. BDA methods are becoming widely adopted also due to the modern computational power enabling Markov Chain Monte Carlo (MCMC) randomized algorithms brooks2011handbook , tools like Stan carpenter2017stan , and libraries like rethinking mcelreath2020statistical and brms burkner2017brms .
@@
    The causal framework entailed by BDA makes causal assumptions explicit. The Bayesian workflow gelman2020bayesian makes any hypothesis of causal relations explicit. Analyses become more transparent, and competing causal assumptions are easier to assess.
@@ 3 Method
We conducted a controlled experiment that investigates the impact of requirements quality on a software engineering activity. Our goal is both to (1) contribute empirical evidence to the effect of quality defects and (2) compare the inferential capabilities of frequentist (FDA) with Bayesian (BDA) statistics. Part of our experiment contributes a conceptual replication baldassarre2014replication of the study conducted and reported by Femmer et al. femmer2014impact and re-analyzed by us frattini2024second , as a subset of our hypotheses overlaps with theirs and our study contributes diagnostic evidence for their claims nosek2020replication . Therefore, we report the design of the experiment with emphasis on the replication following the guidelines by Carver carver2010towards .
@@ 3.1 Goals
RQ1 is dedicated to the main relationship of interest between quality defects and an affected activity. RQ1.1 aligns with the research question driving the original study femmer2014impact , which makes this part of our study a conceptual replication. RQ1.2 extends the scope of the investigation of quality factors with ambiguous pronouns. RQ1.3 investigates the interaction between the two quality factors. RQ2 adds a context-sensitive perspective to the relationship. RQ2.1 focuses on the direct effect that context factors have on the affected activity. RQ2.2 investigates whether context factors mediate the effect of quality defects on the activity.
@@ 3.2 Original Experiment
For the dependent variable, the authors calculated the number of missing domain model elements (i.e., actors, objects, and associations). Although the authors also recorded context variables such as a categorical assessment of general knowledge in SE and RE, these were not used in the analysis. The analysis followed a frequentist approach performing a null-hypothesis significance test for each of the three domain model elements to determine whether a statistically significant difference between the experimental groups exists. The study shows a statistically significant difference in the number of identified associations but not in the number of actors or objects. The authors conclude that the commonly assumed impact of passive voice on missing domain model actors is actually negligible, but passive voice impedes the understanding of the relationships between entities in the requirements specification.
3.3 Reanalysis
The original study by Femmer et al. analyzed its data under simplified assumptions. Among these is the assumption that the three dependent variables (number of missing actors, objects, and associations) only depend on the main factor (use of active or passive voice). We challenged this assumption in a re-analysis of the original data frattini2024second for the following reasons:
  1. 1.
    In a small-scale experiment employing a parallel design, there is no measure to control subject variability vegas2015crossover , such that context factors like experience or skill might affect the dependent variables.
  2. 2.
    Missing an actor or object in the domain model (i.e., a node) necessarily causes an association to be missed (i.e., an edge that would have connected these nodes).
Figure 4(a) visualizes the causal assumptions of the original experiment femmer2014impact as a directed acyclic graph elwert2013graphical (the syntax of which is further explained in Section 3.4.10) and Figure 4(b) shows the revision in scope of the reanalysis frattini2024second . The revision includes (1) two context factors that were already recorded but not used in the original experiment, and (2) two causal relations between the response variables.
(a) Original causal assumptions femmer2014impact
(b) Revised causal assumptions frattini2024second
Figure 4: Causal assumptions about the impact of passive voice
We performed a re-analysis, i.e., an independent analysis of the same data using a different statistical model gomez2010replications , which is sometimes referred to as a test of robustness nosek2020replication . During this re-analysis, we replaced the NHSTs with regression models that include context factors and the affecting response variables in the case of missing associations.
The results of this re-analysis agree with the original study in that the effect of passive voice on the number of missing actors and objects is negligible. However, the re-analysis disagrees with the original study regarding the effect of passive voice on the number of missing associations. The re-analysis determined that passive voice slightly increases the number of missing associations ($\beta_{pv}=0.7$). Still, the confidence interval of this effect ($CI_{pv}=(-0.56,1.90)$) intersects 0 and is, therefore, not significant. On the other hand, the effect of missing objects on missing associations was significant ($\beta_{act}=1.12$, $CI_{act}=(0.32,1.96)$). The re-analysis did not find a significant effect of the available context variables on the response variables. Our re-analysis concludes that the effect of passive voice on the domain modeling activity is less significant than originally assumed frattini2024second .
@@ 3.4 Our Experiment
The reanalysis frattini2024second of the study by Femmer et al. femmer2014impact did improve the conclusion validity of the results but failed to address other shortcomings. For example, the subject variability still threatened the internal validity of the results due to the parallel design of the experiment vegas2015crossover , and the context factors were limited to those recorded during the original study. Hence, we used their study as inspiration for our own presented in this paper and aimed to improve upon the research design. During the preparation of our study, we conferred with the authors of the original study and made the following changes to the original study.
@@
  • Dependent variables: We merged two types of elements in the domain model (the nodes of the model, i.e., actors and objects) into a single type entity because they represent the same concept in the domain model (nodes) femmer2014impact and the distribution in our experimental objects is heavily skewed (16/17 entities are objects). Furthermore, we increased the dependent variables by additionally evaluating the number of superfluous entities, the number of wrong associations, and the duration for creating the domain model.
@@
  • Analysis: The crossover design produces paired data as opposed to the unpaired data of the original experiment, which changes the appropriate hypothesis test vegas2015crossover (Mann-Whitney U test in the original study vs. Wilcoxon signed-rank test in this study). In addition, we extend the original FDA by performing a Bonferroni correction to deal with family-wise error rate when testing multiple hypotheses benjamini1995controlling . Furthermore, we additionally analyze the data using BDA.
Our experiment differs from the original experiment femmer2014impact in all elements gomez2010replications . However, because a subset of our hypotheses aligns with their hypotheses, part of our study counts as a conceptual replication baldassarre2014replication since it contributes diagnostic evidence for the original claims nosek2020replication . In the rest of this subsection, we report the design of our experiment following the guidelines by Jedlitschka et al. jedlitschka2008reporting .
@@ 3.4.2 Hypotheses
The three independent variables $ind\in\{PV,AP,\text{{PVAP}}\}$ (passive voice, ambiguous pronoun, and the coexistence of passive voice and ambiguous pronoun) and the five dependent variables $dep\in\{D,E^{-},E^{+},A^{-},A^{\times}\}$ (duration, missing entities, superfluous entities, missing associations, wrong associations) define our 15 null hypotheses as follows.
$\sum_{ind\in\{PV,AP,\text{{PVAP}}\}}\sum_{dep\in\{D,E^{-},E^{+},A^{-},A^{\times}\}}H_{0}^{ind\rightarrow dep}$
@@
Requirements Quality Defect RQD ind The use of a verb in passive voice, an ambiguous pronoun, or both categorical {none, PV, AP, PVAP}
Experience in SE exp.se con Years of experience in software engineering count $\mathbb{N}$
Experience in RE exp.re con Years of experience in requirements engineering count $\mathbb{N}$
Education edu con Highest acquired degree ordinal {High School, B.Sc., M.Sc., Ph.D.}
Primary role role con SE-related Role with the most years of professional experience categorical {requirements engineer, product owner, software architect, developer, tester, quality engineer, trainer, manager, other, none}
Task experience exp.task con Experience with the task of domain modeling ordinal {never, rarely, from time to time, often}
Formal modeling training formal con Formal training in domain modeling categorical {true,false}
@@
Tool experience tool con Experience with using Google Docs for modeling ordinal {none, rarely, from time to time, often}
Duration $D$ dep Number of minutes it took the participant to complete the experimental task on one requirement count $\mathbb{N}$
Missing entities $E^{-}$ dep Number of relevant entities missing from the submitted domain model count $[0,E_{expected}]$
Superfluous entities $E^{+}$ dep Number of not relevant entities included in the submitted domain model count $\mathbb{N}$
Missing associations $A^{-}$ dep Number of associations missing from the submitted domain model count $[0,A_{expected}]$
Wrong associations $A^{\times}$ dep Number of associations where either the source or target of the edge is a different entity than implied by the requirement count $[0,A_{found}]$
The variables in Table 1 do not include a participant type that distinguishes students from practitioners. While including such a variable is common practice in SE research bogner2023restful , meta-research on the eligibility of students as experiment participants suggests that the labels student or practitioner are merely a proxy for levels of more meaningful factors like domain knowledge and experience salman2015students . Additionally, the line between students and practitioners becomes increasingly blurred as students more commonly gather industrial experience before or during their studies carver2004issues . Consequently, we subsume the participant type variable by the causally more meaningful and fine-grained variables of experience, education, domain knowledge, and formal modeling training. We compared two models—one using the binary distinction and one using the more fine-grained variables—and determined that the latter outperforms the former in predictive power, even though only slightly. This confirms to us that the variables we used are at least as expressive as the binary participant type variable.
@@ 3.4.3 Experimental Design
Given our sampling strategy involving industry practitioners, which are difficult to recruit for controlled experiments sjoberg2003challenges , we anticipated a moderate sample size of participants. Consequently, we opted for a crossover design vegas2015crossover instead of a parallel design, i.e., we apply every treatment to all subjects instead of distributing the subjects among the treatments. Previously, Kitchenham et al. advised against the use of crossover designs kitchenham2003case . Mainly, the validity of crossover design experiments is challenged by the following confounding factors:
@@
    the period in which a treatment is applied to a subject, as certain periods may influence the dependent variables (e.g., participants may mature and perform increasingly better the more often they perform the experimental task subsequently);
@@
    the sequence in which the treatments are applied to a subject, as certain sequences may have a beneficial effect on a dependent variable (e.g., there might be an optimal sequence to apply the treatments in);
@@
    the effect from a previous treatment may carry over to the period when applying a subsequent treatment vegas2015crossover ; and
  4. 4.
    the subject variability, as software engineering tasks are highly dependent on the skill of involved individuals pickard1998combining .
However, recent adoptions of best practices from other disciplines made this design applicable to SE research without compromising the validity of the results vegas2015crossover ; fucci2016an . The threats to validity can be mitigated during design and analysis vegas2015crossover by (1) randomizing the order of treatments and (2) including the independent variables’ period, sequence, the interaction between them (representing the carryover effect), and subject variability in the analysis. Consequently, we consider the experimental design variables listed in Table 2 in addition to the variables listed in Table 1.
Table 2: Variables of the study (experimental design factors)
Variable Description Data type Range
Period Index of the experimental period in which the data was obtained ordinal [1; 4]
Sequence Order in which a participants received the treatments categorical {1234, 1243, …, 4321}
Carryover effect Interaction between the period and the treatment categorical $1\times none$, $1\times PV$, …, $4\times PVAP$
Subject variability Index of a participant categorical {1, 2, …, 25}
When controlling the threats to validity, the crossover design provides two benefits. Firstly, it requires fewer participants, as an experiment with $n_{p}$ participants and $n_{t}$ treatments yields $n_{p}\times n_{t}$ observations instead of only $n_{p}$ wohlin2012experimentation . Secondly, it accounts for subject variability, as the dependent variables can be measured in relation to the average response of each subject instead of the average response of each treatment group kitchenham2003case . Therefore, each subject acts as its own control and mitigates within-subject variability.